x is directly proportional to y^2. When y is 3, x is 6. Then y is 6, what is x?
A) 6 B) 10 C) 12 D) 18 E) 24

Answer by susanaramya(12)

x is propotional to y^2
x=k y^2,k is a constant
when y=3,x=6 6=(3^2)k
k=6/9=2/3
the equation is x=2(y^2)/3
when y=6 x=2(6^2)/3
x=2(36)/3
x=24
so answere is (E)24